OpenStudy (anonymous):

48x56=3x7x2c what is the value of c

OpenStudy (anonymous):

45 x 56 = 3x7x2c 2520= 42c c= 60

OpenStudy (anonymous):

c = 48 * 56/ 3 * 7 * 2 = 16 * 4 = 64